# Scalable Graph Condensation with Evolving Capabilities

### v1

@@ Abstract.
Graph data has become a pivotal modality due to its unique ability to model relational datasets. However, real-world graph data continues to grow exponentially, resulting in a quadratic increase in the complexity of most graph algorithms as graph sizes expand. Although graph condensation (GC) methods have been proposed to address these scalability issues, existing approaches often treat the training set as static, overlooking the evolving nature of real-world graph data. This limitation leads to inefficiencies when condensing growing training sets. In this paper, we introduce GECC (Graph Evolving Clustering Condensation), a scalable graph condensation method designed to handle large-scale and evolving graph data. GECC employs a traceable and efficient approach by performing class-wise clustering on aggregated features. Furthermore, it can inherits previous condensation results as clustering centroids when the condensed graph expands, thereby attaining an evolving capability. This methodology is supported by robust theoretical foundations and demonstrates superior empirical performance. Comprehensive experiments show that GECC achieves better performance than most state-of-the-art graph condensation methods while delivering an around 1,000$\times$ speedup on large datasets.
@@ 1. Introduction
Graph-structured data has become indispensable in various domains, including social networks (Fan et al., 2019), epidemiology (Liu et al., 2024), and recommendation systems (Wu et al., 2020; Wang et al., 2024b). The ability of graphs to represent complex relationships and dependencies has propelled their adoption in machine learning, especially with the advent of graph neural networks (GNNs) (Bronstein et al., 2017). However, the exponential growth of real-world graph datasets presents significant computational challenges, as the cost of training GNNs increases with the number of nodes and edges (Hu et al., 2021; Huang et al., 2021). To address this, graph condensation (GC) techniques (Jin et al., 2022b, a; Wang et al., 2024a; Zhang et al., 2024; Zheng et al., 2024a; Xiao et al., 2024; Liu and Shen, 2024) have been developed, which aim to produce significantly smaller yet information-rich graphs that accelerate GNN training. For example, GCond (Jin et al., 2022b) condenses the Flickr dataset to 0.1% of its original size while preserving 99.8% of the original accuracy. These condensed graphs not only save significant storage space and transmission bandwidth, but also enhance the efficiency of retraining neural networks in many critical applications such as continual learning (Liu et al., 2023).
Table 1. Condensation time on evolving graphs, reported as a multiple of GNN training time (21.37s) on uncondensed graphs (Reddit) in the final time step.
@@
GCond 75× 112× 149× 104× 185×
GCond-Init 75× 89× 107× 120× 134×
GCond-Grow 75× 94× 112× 101× 119×
@@
Despite the promise, existing GC methods face three significant limitations that hinder their practical utility in real-world scenarios.
@@
(a) High Computational Overhead: Most GC methods entail a computationally intensive condensation process, marked by frequent gradient calculations and updates. This process closely resembles full GNN training, creating a paradox as it contradicts the primary objective of GC methods to enhance the efficiency of GNN training. For instance, gradient matching techniques (Jin et al., 2022b, a) require consistent alignment of GNN gradients with each iteration. Similarly, trajectory matching methods (Zheng et al., 2024a; Zhang et al., 2024) necessitate the alignment of GNN parameters at various stages during the training process. These methods inherently demand significant GNN training and gradient processing, thus slowing down the procedure as the size of graph datasets expands.
(b) Challenges with Evolving Graphs: Existing GC methods are designed for static graphs and overlook the evolving nature of real-world graphs. In practice, graphs evolve over time, with nodes, edges, or attributes being added, removed, or modified. Existing graph condensation (GC) methods require re-running the entire condensation process from scratch whenever the training set changes, since modifications affect every synthetic node. Moreover, while synthetic graphs are expected to grow proportionally with the original graphs, an effective strategy for accommodating this growth is currently lacking.
@@
In Table 1, we report the relative condensation times, i.e., multiple of a standard GNN training time, for evolving graphs on the Reddit dataset, using GCond (Jin et al., 2022b) as a representative example. We compare three variants: (i) GCond: Performs condensation from scratch at each time step. (ii) GCondX-Init: Retains synthetic nodes from the previous time step as initialization, updating all nodes during the condensation process. (iii) GCond-Grow: Freezes the synthetic nodes from the previous time step and concatenate new synthetic nodes to grow the condensed graph. GCond-Init demonstrates the extent to which evolving adaptation strategies can accelerate the convergence speed of GC. Meanwhile, GCond-Grow illustrates how directly reducing complexity enhances the overall efficiency of GC. The table indicates that the condensation process requires over 100$\times$ more time than the GNN training on the uncondensed graph. This substantial computational overhead suggests that current GC methods may not be practical for evolving graphs. Even though simple evolving strategies can reduce marginal time, the cost remains significantly higher than that of GNN training, underscoring the urgent need for a novel GC method that can efficiently handle graph evolution.
@@
(c) Lack of Traceability: Many GC methods synthesize condensed graphs without explicitly establishing a connection between the nodes in the condensed graph and those in the original graph. This lack of explicit connections obscures the contribution of each original node to its synthetic counterpart, diminishing the traceability of the condensation process. Although these condensed nodes might be informative to GNN models, discerning their real-world semantic meanings can be challenging to humans. This poor traceability can significantly limit the application of GC in sectors where clarity, comprehensive data explanations, and transparent decision-making are crucial. Also, with traceability, users can manipulate the condensed graph alongside the original one—for example, easily filtering low-quality data when the corresponding original nodes exhibit poor quality.
@@
To tackle the three critical challenges altogether, we introduce a novel framework called Graph Evolving Clustering Condensation (GECC). We begin by refining our objectives for GC and providing theoretical bounds that elucidate the connection between GC and clustering. This insight allows us to perform condensation efficiently without resorting to the expensive gradient computations typically required for training GNNs on the entire dataset or other learnable modules. Specifically, we cluster propagated node features into partitions, where centroids serve as the condensed node features at each time step. Additionally, this method adapts well to evolving, expanding graphs through incremental clustering, using centroids from previous time steps as initializations for subsequent steps. Notably, GECC enhances traceability by establishing a clear correspondence between condensed and original nodes, offering deeper insights into how condensed nodes encapsulate information from the actual graph. Our contributions are outlined as follows:
@@
  2. Leveraging our theoretical insights, we introduce the first training-free framework designed to significantly accelerate condensation and adapt to the evolving nature of real-world graphs, while providing favorable traceability.
  3. Comprehensive experiments demonstrate that GECC achieves state-of-the-art (SOTA) accuracy and efficiency in GC. For example, GECC can condense the evolving Reddit dataset more than 1000 times faster than GCond, while outperforming other existing methods in terms of test accuracy.
@@ 2.1. Model-based Graph Condensation
Model-based GC methods, which require a GNN training phase on the original graph, were the first approaches proposed for GC. In these methods, a smaller graph is synthesized to effectively represent the original for training GNNs (Hashemi et al., 2024). For example, GCond (Jin et al., 2022b), DosCond (Jin et al., 2022a), and SGDD (Yang et al., 2023) minimize a gradient matching loss that aligns the gradients of the training losses w.r.t. the GNN parameters computed on both the original and condensed graphs. Alternatively, SFGC (Zheng et al., 2024b) and GEOM (Zhang et al., 2024) condense a graph by aligning the parameter trajectories of the original graph, thereby eliminating the need for explicit edge generation and reducing complexity. Although these approaches are effective, they require significant computational resources due to the multiple full GNN training runs on the original graph. To mitigate this cost, GCDM (Liu et al., 2022) and SimGC (Xiao et al., 2024) adopt a distribution matching strategy by aligning sampled original node features with synthetic nodes, thus reducing the discrepancy between gradient matrices and the final output. Other methods such as GCSNTK (Wang et al., 2024a) use GNTK (Du et al., 2019) to bypass the gradient update in the synthetic graph; however, such approaches currently show suboptimal performance according to recent benchmarks (Gong et al., 2024; Sun et al., 2024). Overall, these model-based GC methods share common challenges including high computational overhead, scalability, and traceability issues. Recent work MCond (Gao et al., 2023) explicitly learns a one-to-many node mapping from original nodes to synthetic nodes, thereby facilitating inductive representation learning and enhancing traceability; however, it still operates within the framework of gradient matching and remains highly complex.
@@ 2.2. Model-Agnostic Graph Condensation
While model-based approaches have advanced GC, they still struggle to scale to large-scale graph datasets (e.g., graphs with over one million nodes such as Ogbn-products (Hu et al., 2021)). To further reduce condensation time on such large-scale graphs, recent efforts have focused on directly matching the representations between the original and synthetic graphs, bypassing the need for a surrogate model.. This strategy offers benefits in efficiency and generalizability to diverse downstream models. GCPA (Li et al., 2025) first extracts structural and semantic information from the original graph, randomly samples the representations, and finally refines them using contrastive learning. Likewise, CGC (Gao et al., 2024) revisits the node feature matching paradigm by reducing the representation matching problem to a class partitioning task. In this approach, nodes within the same partition are merged, and a learning module is employed to weight different samples before they are aggregated into a synthetic node. However, despite its effectiveness, CGC’s theoretical analysis does not establish a concrete relationship between class partitioning and the GC objective, thereby questioning the necessity of this extra step. Moreover, these model-agnostic methods still require an additional learning module, which introduces challenges in optimization and hyperparameter tuning. Finally, none of these methods fully address the evolving nature of large-scale graphs, which typically necessitates a complete re-condensation whenever the graph changes. To our best knowledge, GECC is the first model-agnostic and training-free method with linear complexity.
@@ 3. Preliminaries and Notations
Given a node set $\mathcal{V}$ and an edge set $\mathcal{E}$, a graph is denoted as $G=({\mathcal{V},\mathcal{E}})$. In the case of attributed graphs, where nodes are associated with features, the graph can be represented as $G=({\bf{X},\bf{A}})$, where ${\bf X}=[{\bf x}_{1},{\bf x}_{2},...,{\bf x}_{N}]$ denotes the node attributes, and ${\bf A}$ shows the adjacency matrix. The graph Laplacian matrix is $\bf L=\bf D-\bf A$, where $\bf D$ is a diagonal degree matrix with ${\bf D}_{ii}=\sum_{j}{\bf A}_{ij}$. Let $N=|\mathcal{V}|$ and $E=|\mathcal{E}|$ represent the number of nodes and edges, respectively.
@@
Graph Condensation Formulation. GC aims to condense a smaller synthetic graph $G^{\prime}=(\mathbf{X}^{\prime},\mathbf{A}^{\prime})$, where $\mathbf{X}^{\prime}\in\mathbb{R}^{N^{\prime}\times d}$, $\mathbf{A}^{\prime}\in\{0,1\}^{N^{\prime}\times N^{\prime}}$, and $N^{\prime}\ll N$, from the original large graph $G=(\mathbf{X},\mathbf{A})$. The objective is to ensure that GNNs trained on $G^{\prime}$ achieve performance comparable to those trained on $G$, thereby significantly accelerating GNN training (Jin et al., 2022b). The large-scale graph $G_{t}=(\mathbf{X}_{t},\mathbf{A}_{t})$ serves as the original graph of our GC framework. Each node is associated with one of $c$ classes, encoded as numeric labels $\mathbf{y}_{t}\in\{1,\dots,c\}^{N_{t}}$ and one-hot labels $\mathbf{Y}_{t}\in\mathbb{R}^{N_{t}\times c}$. Any GC method focuses on generating a condensed graph $G^{\prime}_{t}=(\mathbf{X}^{\prime}_{t},\mathbf{A}^{\prime}_{t})$ from the original graph $G_{t}=(\mathbf{X}_{t},\mathbf{A}_{t})$, preserving the key structural and feature information required for downstream tasks.
@@
Evolving Graph Condensation Formulation. In the evolving graph scenario, we consider a sequential stream of graph batches $\{B_{1},B_{2},\dots,B_{m}\}$, where $m$ represents the total number of time steps. Each graph batch $B_{i}=(\mathbf{X}_{i},\mathbf{A}_{i})$ contains newly added nodes along with their associated edges in inductive graphs, while in transductive graphs, it contains newly labeled nodes but retains the entire graph structure (Su et al., 2023). These graph batches are progressively integrated into the existing graph over time, constructing a series of incremental graphs $\{G_{1},G_{2},\dots,G_{t}\}$. At time step $t$, the snapshot graph $G_{t}=(\mathbf{X}_{t},\mathbf{A}_{t})$ encompasses all nodes and edges that have appeared up to that point, with $G_{t}=\bigcup_{i=1}^{t}B_{i}$, and importantly, we preserve the distribution of classes across splits as the graph evolves. At each time step $t$, during the GC phase, we aim to generate a condensed graph $G^{\prime}_{t}=(\mathbf{X}^{\prime}_{t},\mathbf{A}^{\prime}_{t})$ from the snapshot graph $G_{t}$. This condensed graph $G^{\prime}_{t}$, which retains the essential structural and feature information of $G_{t}$, is then used to train GNNs efficiently for any downstream tasks. During deployment, the GNN trained on $G^{\prime}_{t}$ is applied to classify the nodes in $G_{t}$. As new graph batches are added at time step $t+1$, the graph expands to $G_{t+1}$. Unlike previous approaches that require repeating the GC process from scratch, our method can effectively inherit the condensed graph from the previous time step with minimal computational overhead, ensuring that $G^{\prime}_{t+1}$ effectively represents the expanded graph $G_{t+1}$ while maintaining high performance on the growing dataset.

### v2

@@ Abstract.
The rapid growth of graph data creates significant scalability challenges as most graph algorithms scale quadratically with size. To mitigate these issues, Graph Condensation (GC) methods have been proposed to learn a small graph from a larger one, accelerating downstream tasks. However, existing approaches critically assume a static training set, which conflicts with the inherently dynamic and evolving nature of real-world graph data. This work introduces a novel framework for continual graph condensation, enabling efficient updates to the distilled graph that handle data streams without requiring costly retraining. This limitation leads to inefficiencies when condensing growing training sets. In this paper, we introduce GECC (Graph Evolving Clustering Condensation), a scalable graph condensation method designed to handle large-scale and evolving graph data. GECC employs a traceable and efficient approach by performing class-wise clustering on aggregated features. Furthermore, it can inherit previous condensation results as clustering centroids when the condensed graph expands, thereby attaining an evolving capability. This methodology is supported by robust theoretical foundations and demonstrates superior empirical performance. Comprehensive experiments including real world scenario show that GECC achieves better performance than most state-of-the-art graph condensation methods while delivering an around 1000$\times$ speedup on large datasets.
@@ 1. Introduction
Graph-structured data has become indispensable in various domains, including social networks (Fan et al., 2019), epidemiology (Liu et al., 2024), and recommendation systems (Wu et al., 2020). The ability of graphs to represent complex relationships and dependencies has propelled their adoption in machine learning, especially with the advent of graph neural networks (GNNs) (Bronstein et al., 2017). However, the exponential growth of real-world graph datasets presents significant computational challenges, as the cost of training GNNs increases with the number of nodes and edges (Hu et al., 2021; Huang et al., 2021). To address this, graph condensation (GC) techniques (Jin et al., 2021, 2022; Wang et al., 2024; Zhang et al., 2024; Zheng et al., 2024; Xiao et al., 2024) have been developed, which aim to produce significantly smaller yet information-rich graphs that accelerate GNN training. For example, GCond (Jin et al., 2021) condenses the Flickr dataset to 0.1% of its original size while preserving 99.8% of the original accuracy on node classification. These condensed graphs not only save significant storage space and transmission bandwidth, but also enhance the efficiency of retraining neural networks in many critical applications such as continual learning (Liu et al., 2023).
Table 1. Condensation time on evolving graphs, reported as a multiple of GNN training time (21.37 s) on uncondensed graphs (Reddit, 232,965 nodes) in the final time step. The GNN training time on condensed graphs is 1.1 s.
@@
GCondX 75× 112× 149× 104× 185×
GCondX-Init 75× 89× 107× 120× 134×
GCondX-Grow 75× 94× 112× 101× 119×
@@
Despite the promise, existing GC methods face three significant limitations that limit their practical utility in real-world scenarios.
@@
(a) High Computational Overhead: Most GC methods entail a computationally intensive condensation process, marked by repeated gradient calculations and updates. This process closely resembles full GNN training, creating a paradox as it contradicts the primary objective of GC methods to enhance the efficiency of GNN training. For instance, gradient matching techniques (Jin et al., 2021, 2022) require consistent alignment of GNN gradients across successive iterations. Similarly, trajectory matching methods (Zheng et al., 2024; Zhang et al., 2024) necessitate the alignment of GNN parameters at various stages during the training process. These methods inherently demand significant GNN training and gradient processing, thus slowing down the procedure as the size of graph datasets expands.
@@
(b) Challenges with Evolving Graphs: Existing GC methods are designed for static graphs and overlook the evolving nature of real-world graphs. In practice, graphs evolve over time, with nodes, edges, or attributes being added, removed, or modified. Existing graph condensation (GC) methods require re-running the entire condensation process from scratch whenever the training set changes, since modifications affect every synthetic node. Moreover, while synthetic graphs are expected to grow proportionally with the original graphs, an effective strategy for adapting to this growth is currently lacking. In Table 1, we report the relative condensation times for evolving graphs on the Reddit dataset, using GCondX (Jin et al., 2021) as a representative example. We compare three variants: (i) GCondX: Performs condensation from scratch at each time step. (ii) GCondX-Init: Retains synthetic nodes from the previous time step as initialization, updating all nodes during the condensation process. (iii) GCondX-Grow: Freezes the synthetic nodes from the previous time step and concatenate new synthetic nodes to grow the condensed graph. The table indicates that the condensation process requires over 100$\times$ more time than the GNN training on the uncondensed graph. This substantial computational overhead suggests that current GC methods may not be practical for evolving graphs. Even though simple evolving strategies can reduce marginal time, the cost remains significantly higher than that of GNN training, underscoring the urgent need for a novel GC method that can efficiently handle graph evolution.
@@
(c) Lack of Traceability: Many GC methods synthesize condensed graphs without explicitly establishing a connection between the nodes in the condensed graph and those in the original graph. This lack of explicit connections obscures the contribution of each original node to each condensed node, diminishing the traceability of the condensation process. Although these condensed nodes might be informative to GNN models, discerning their real-world semantic meanings can be challenging to humans. This poor traceability can significantly limit the application of GC in sectors where clarity, comprehensive data explanations, and transparent decision-making are crucial. Also, with traceability, users can manipulate the condensed graph alongside the original one—for example, easily filtering low-quality data when the corresponding original nodes exhibit poor quality.
@@
To tackle the three critical challenges all together, we introduce a novel framework called Graph Evolving Clustering Condensation (GECC). We begin by refining our objectives for graph condensation and establishing theoretical upper bounds that clarify the connection between graph condensation and clustering. This connection enables us to perform condensation efficiently without the costly gradient calculations needed for traditional GNN training. Specifically, we cluster propagated node features into partitions, where centroids serve as the condensed node features at each time step. Additionally, this method adapts well to evolving, expanding graphs through incremental clustering, using centroids from previous time steps as initializations for subsequent steps. Notably, GECC enhances traceability by establishing a clear correspondence between condensed and original nodes, offering deeper insights into how condensed nodes encapsulate information from the actual graph. Our contributions are outlined as follows:
@@
  2. Leveraging our theoretical insights, we introduce the first framework designed to significantly accelerate condensation and adapt to the evolving nature of real-world graphs, while providing traceability. This framework methodically partitions node representations to produce cluster centroids at each time step, which serve as the condensed node features.
  3. Comprehensive experiments demonstrate that GECC achieves state-of-the-art (SOTA) accuracy and efficiency in GC. For example, GECC can condense the evolving Reddit dataset more than 1000 times faster than GCond, while outperforming other existing methods in terms of accuracy.
@@ 2.1. Model-based Graph Condensation
Model-based GC methods, which require a GNN training phase on the original graph, were the first approaches proposed for GC. In these methods, a smaller graph is synthesized to effectively represent the original for training GNNs (Hashemi et al., 2024). For example, GCond (Jin et al., 2021), DosCond (Jin et al., 2022), and SGDD (Yang et al., 2023) minimize a gradient matching loss that aligns the gradients of the training losses w.r.t. the GNN parameters computed on both the original and condensed graphs. Alternatively, SFGC (Zheng et al., 2024) and GEOM (Zhang et al., 2024) condense a graph by aligning the parameter trajectories of the original graph, thereby eliminating the need for explicit edge generation and reducing complexity. Although these approaches are effective, they require significant computational resources due to the multiple full GNN training runs on the original graph. To mitigate this cost, GCDM (Liu et al., 2022) and SimGC (Xiao et al., 2024) adopt a distribution matching strategy by aligning sampled original node features with synthetic nodes, thus reducing the discrepancy between gradient matrices and the final output. Other methods such as GCSNTK (Wang et al., 2024) use GNTK (Du et al., 2019) to bypass the gradient update in the synthetic graph; however, such approaches currently show suboptimal performance according to recent benchmarks (Gong et al., 2024; Sun et al., 2024). OpenGC (Gao et al., 2024b) enhances graph condensation for evolving graphs by simulating structure-aware distribution shifts to extract temporal invariant patterns, thereby improving the generalization of GNNs in real-world scenarios. Overall, these model-based GC methods share common challenges including high computational overhead, scalability, and traceability issues. The recently proposed MCond (Gao et al., 2024a) explicitly learns a one-to-many node mapping from original nodes to synthetic nodes, thereby facilitating inductive representation learning and enhancing traceability; however, it still operates within the framework of gradient matching and remains highly complex.
@@ 2.2. Model-Agnostic Graph Condensation
While model-based approaches have advanced GC, they still struggle to scale to large-scale graph datasets (e.g., graphs with over one million nodes such as Ogbn-products (Hu et al., 2021)). To further reduce condensation time on such large-scale graphs, recent efforts have focused on directly matching the representations between the original and synthetic graphs, bypassing the need for a surrogate model, i.e., the GNN model used during the condensation. This strategy offers benefits in efficiency and generalizability to diverse downstream models. GCPA (Li et al., 2025) first extracts structural and semantic information from the original graph, randomly samples the representations, and finally refines them using contrastive learning. Likewise, CGC (Gao et al., 2024c) reduces the representation matching problem to a class partitioning task. In this approach, nodes within the same partition are merged, and a learning module is employed to weight different samples before they are aggregated into a synthetic node. However, despite its effectiveness, CGC’s theoretical analysis does not establish a relationship between class partitioning and the GC objective, thereby questioning the necessity of this extra step. Moreover, these model-agnostic methods still require a learning module, which introduces challenges in optimization and hyperparameter tuning. Finally, none of these methods fully address the evolving nature of large-scale graphs, which typically necessitates a complete re-condensation whenever the graph changes. To our best knowledge, GECC is the first model-agnostic and training-free method with linear complexity.
@@ 3. Preliminaries and Notations
Given a node set $\mathcal{V}$ and an edge set $\mathcal{E}$, a graph is denoted as $G=({\mathcal{V},\mathcal{E}})$. In the case of attributed graphs, where nodes are associated with features, the graph can be represented as $G=({\bf{X},\bf{A}})$, where ${\bf X}=[{\bf x}_{1},{\bf x}_{2},...,{\bf x}_{N}]$ denotes the node attributes, and ${\bf A}$ denotes the adjacency matrix. The graph Laplacian matrix is $\bf L=\bf D-\bf A$, where $\bf D$ is a diagonal degree matrix with ${\bf D}_{ii}=\sum_{j}{\bf A}_{ij}$. Let $N=|\mathcal{V}|$ and $E=|\mathcal{E}|$ represent the number of nodes and edges, respectively.
3.1. Graph Condensation.
@@
GC aims to create a smaller synthetic graph $G^{\prime}=(\mathbf{X}^{\prime},\mathbf{A}^{\prime})$, where $\mathbf{X}^{\prime}\in\mathbb{R}^{N^{\prime}\times d}$, $\mathbf{A}^{\prime}\in\{0,1\}^{N^{\prime}\times N^{\prime}}$, and $N^{\prime}\ll N$, from the original large graph $G=(\mathbf{X},\mathbf{A})$. The objective is to ensure that GNNs trained on $G^{\prime}$ achieve performance comparable to those trained on $G$, thereby significantly accelerating GNN training (Jin et al., 2021). The large-scale graph $G=(\mathbf{X},\mathbf{A})$ serves as the original graph of our GC framework. Each node is annotated with one of $c$ classes, encoded as numeric labels $\mathbf{y}\in\{1,\dots,c\}^{N}$ and one-hot labels $\mathbf{Y}\in\mathbb{R}^{N\times c}$. Any GC method focuses on finding an update function $\mathcal{F}$ that generates a new condensed graph $G^{\prime}=\mathcal{F}(G)$, preserving the key structural and feature information required for downstream tasks.
3.2. Evolving Graph Condensation.
Motivation: Training GNNs on real-world graphs is challenging because these graphs constantly evolve, leading to significant computational costs. Current GC techniques, which create smaller training graphs, are designed for static snapshots. Consequently, adapting to graph evolution requires re-running the entire expensive condensation process at each step. This wasteful approach motivates our work on Evolving Graph Condensation, which aims to efficiently update a condensed graph incrementally rather than rebuilding it from scratch. In the evolving graph scenario, we consider a sequential stream of graph batches $\{B_{1},B_{2},\dots,B_{m}\}$, where $m$ represents the total number of time steps. Each graph batch $B_{i}=(\mathbf{X}_{i},\mathbf{A}_{i})$ contains newly added nodes along with their associated edges in inductive graphs, while in transductive graphs, it contains newly labeled nodes but retains the entire graph structure (Su et al., 2023).
@@
Problem Formulation: In the evolving graph scenario, we consider a sequential stream of graph batches $\{B_{1},B_{2},\dots,B_{m}\}$, where $m$ is the total number of time steps. Each graph batch $B_{i}=(\mathbf{X}_{i},\mathbf{A}_{i})$ introduces new information. In inductive graphs, a batch contains newly added nodes and their edges. In transductive graphs, it typically contains newly revealed labels for existing nodes while the overall structure is retained (Su et al., 2023). These batches are progressively integrated, constructing a series of snapshot graphs $\{G_{1},G_{2},\dots,G_{m}\}$. At any time step $t$, the snapshot graph $G_{t}=(\mathbf{X}_{t},\mathbf{A}_{t})$ encompasses all information up to that point, where $G_{t}=\bigcup_{i=1}^{t}B_{i}$. Given the snapshot graph $G_{t}$ at the current time step $t$ and the condensed graph $G^{\prime}_{t-1}$ from the previous step, the goal of evolving graph condensation is to find an update function $\mathcal{F}$ that generates a new condensed graph $G^{\prime}_{t}=\mathcal{F}(G^{\prime}_{t-1},B_{t})$. The objective is for $G^{\prime}_{t}$ to be significantly smaller than $G_{t}$ while preserving essential structural and feature information, such that any GNN trained on $G^{\prime}_{t}$ achieves performance on downstream tasks that is comparable to a GNN trained on the full, original graph $G_{t}$. At each time step $t$, instead of repeating the GC process from scratch on $G_{t}$, our method aims to efficiently update the condensed graph. It effectively inherits knowledge from the previously condensed graph $G^{\prime}_{t-1}$ and integrates the new information from batch $B_{t}$ to produce $G^{\prime}_{t}$. This updated graph is then used to train GNNs, which are deployed to classify nodes in the full graph $G_{t}$. This incremental approach ensures that the condensed representation stays current with the evolving graph while minimizing redundant computation.